One angle of a triangle is 20degrees° greater than the smallest​ angle, and the third angle is 40degrees° less than twice the smallest angle. Find the measures of the three angles.
Two complementary angles are ​(9x−2)degrees° and ​(5x+8​)degrees°. Find their measures.
​(9x−2​)degrees°equals= ​(Simplify your answer. Type an integer or a​ fraction.)
Thank you!

Answer by addingup(3677)

the angle of every triangle add up to 180 degrees.
let the smallest angle be x.
one angle: x+20
a third: 2x-40
------------------------
x+x+20+2x-40 = 180
4x-20 = 180
4x = 200
x = 50
--------------------------
smallest angle: 50
another angle: 50+20 = 70
a third: 2(50)-40 = 60
50+70+60 = 180 Correct